OpenStudy (anonymous):

solve: 3x-5=5x + 2 how to solve this inequality?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3x-5%3D5x+%2B+2

OpenStudy (anonymous):

3x-5=5x+2 -5-2=5x-3x -7/2=x

OpenStudy (anonymous):

it has the answer there

OpenStudy (anonymous):

is it okay

OpenStudy (anonymous):

thanks. :3 but, when we put -7/2 is it possible to input it into the equation as so?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

like i said the answer is there on the web site wolf ram is a good and helpful place

OpenStudy (anonymous):

Thank you guys. ^^;

hero (hero):

Looks like an equality rather than inequality

OpenStudy (anonymous):

its an inquelitey

hero (hero):

I thought inequalities had symbols that included > or <

OpenStudy (anonymous):

oh, my bad it is an equality. xD i must be reading my text book wrong.